OpenStudy (anonymous):

For each of the following questions, choose the correct answer. Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Write a system of equations to model the problem if x represents Joelle’s score and y represents Anita’s score.

OpenStudy (anonymous):

A. x + y = 40 y = 2x – 425 B. x + y = 425 y = 2x + 40 C. x + y = 425 y = 2x – 40 D. x – y = 425 y = 2x – 40

OpenStudy (sleepyjess):

Lets use A for anita and J for Joelle. A=2J-40

OpenStudy (sleepyjess):

A+J=425

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

and how does that help 1497 learn how to do this?

OpenStudy (anonymous):

sleepyjess already gave you answer, just with different variables, A instead of x and J instead of y

OpenStudy (sleepyjess):

Oh, I thought you needed the actual answer.....

OpenStudy (sleepyjess):

Not just the system of equations